# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

### v1

@@ 3.2 FedMKT
Bidirectional Token Alignment with MinED. A notable challenge that arises during the alignment of output logits distributions is the mismatch between the tokenizers of different LLMs and SLMs, such as the Bloom tokenizer and the LLaMa tokenizer. Consider the sentence, ”we utilize the dynamic programming approach to align tokens” as an example. Utilizing the Bloom tokenizer would segment it into the following tokens: [’we’, ’utilize’, ’the,’ ’dynamic,’ ’programming,’ ’approach,’ ’to,’ ’align,’ ’tokens’]. However, if the LLaMa tokenizer were used, the segmentation would be: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. This misalignment, coupled with model heterogeneity, poses a significant hurdle. Therefore, achieving token alignment across multiple SLMs and LLM is paramount for effective knowledge transfer in a federated learning setting. Proper alignment guarantees the accurate mapping of output logits and probabilistic distribution matrices.
To tackle this issue, we utilize dynamic programming techniques that foster resilient alignment and minimize the impact of tokenizer mismatches, as evidenced in studies Wan et al. (2024); Fu et al. (2023). Given that tokens produced by distinct tokenizers for the same sequence often exhibit minimal disparities, we employ a minimum edit distance (MinED) strategy. This approach maps tokens from various tokenizers based on their MinED, allowing for flexible token alignment. By relaxing the alignment criteria, we retain significant information within the distribution matrices while introducing negligible errors. This method ensures effective knowledge transfer in the federated learning setting, overcoming the challenges posed by tokenizer mismatches and model heterogeneity.
Knowledge transfer between LLMs and SLMs involves a bidirectional token alignment process, which is divided into two stages.
  • In the first stage, clients transfer knowledge of their SLMs to the LLM on the server, aiming to enhance the LLM’s capacity. During this phase, the SLMs generate output logits that are uploaded to the server. Here, the tokens produced by the SLMs serve as the source, while the tokens from the LLM serve as the target. It is crucial to accurately align the SLM tokens with the corresponding LLM tokens.
  • In the second stage, the LLM on the server transfers its knowledge back to the clients’ SLMs, aiming to improve their performance. As the LLM’s output logits are downloaded to the clients, the roles reverse: the LLM’s tokens become the source, and the SLM tokens become the target.
This bidirectional alignment ensures seamless integration and effective knowledge transfer between the models. We can address the challenges posed by tokenizer mismatches and model heterogeneity by leveraging this approach.
Selective Mutual Knowledge Transfer Between LLM and SLMs. Each client $k$ inserts a local small low-rank adapter parameterized by $\theta_{k}$ into its local SLM. We denote client $k$ local SLM with the added $\theta_{k}$ as $g_{\phi_{k}+\theta_{k}}$. Each client $k$ then trains its local LoRA adapter module $\theta_{k}$ using its private dataset. Consequently, Eq.(1) can be reformulated as follows:
$\mathcal{L}_{1}(\theta_{1},\theta_{2},...,\theta_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (4)
$=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathcal{D}^{k}}$ $\ell_{\text{TA}}^{k}(g_{\phi_{k}+\theta_{k}}(x),y)$
@@
where $\ell_{\text{TA}}$ is the task loss for training the local LoRA module $\theta_{k}$ in the client $k$. The original model parameter $\phi_{k}$ of client $k$’s local SLM is frozen during training.
The server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$. We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$. The server conducts the knowledge transfer from each local SLM $g_{\phi_{k}+\theta_{k}}$ to the LLM $f_{\psi+\omega}$ and through supervised fine-tuning and selective knowledge distillation based on the public dataset $\mathcal{D}^{p}$.
@@
We formulate the losses of supervised fine-tuning $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
$\mathcal{L}^{f}_{\text{FT}}(\omega;\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(f_{\psi+\omega}(x),y)$ (5)
$\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
@@
The selective mutual knowledge distillation losses for $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ models (denoted as $\mathcal{L}_{\text{KD}}^{f}$ and $\mathcal{L}_{\text{KD}}^{g}$) are formulated as follows:
$\mathcal{L}^{f}_{\text{KD}}(\omega;\mathcal{D}^{p},\theta_{1},\theta_{2},...,\theta_{K})$ (7)
$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(f_{\psi+\omega}(x),g_{\phi_{k}+\theta_{k}}(x))$
$\mathcal{L}^{g}_{\text{KD}}(\theta_{k};\mathcal{D}^{p},\psi+\omega)$ (8)
$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),f_{\psi+\omega}(x))$
Since clients’ knowledge may adversely affect LLM on the server side in heterogeneous environments, and the reverse is also true, where the server-side LLM could potentially have a negative influence on the clients’ SLMs, it is essential to evaluate the prediction quality of both LLM and SLMs for each public data sample. To facilitate positive knowledge transfer, we propose selective strategies on both the server and client sides, termed DualMinCE. DualMinCE work as follows: For each public sample, we compute the cross-entropy loss across all SLMs and LLM, comparing the prediction distribution with the gold labels. Subsequently, when knowledge needs to be transferred from SLMs to the LLM, we select the SLM exhibiting the minimum cross-entropy loss to effectively distill its knowledge into the LLM residing on the server. Conversely, when transferring knowledge from the LLM to SLMs, we only proceed with the distillation if the LLM’s cross-entropy loss is lower than that of the SLMs. This ensures that the most accurate and informative knowledge is distilled between the LLM and SLMs.
@@
where $\lambda$ is the hyperparameter that control the weight of mutual knowledge transfer.
FedMKT Workflow. FedMKT cultivates a mutually beneficial relationship between the server and clients, fostering a symbiosis where both parties reap the rewards of collective knowledge and expertise embedded in their respective language models. By harnessing the distinct strengths of server-side LLM and client-side SLMs, FedMKT pioneers a new frontier in efficient and effective federated learning within the realm of LLM. This innovative approach enables the collaborative evolution of LLM and SLMs. The workflow of FedMKT is outlined as follows:
  1. 1.
  2. 2.
    After local training, the $K$ clients upload their SLMs’ output logits and the corresponding cross-entropy losses computed on the public dataset to the server.
  3. 3.
  4. 4.
    On the server side, knowledge is seamlessly selectively transferred from the SLMs to the LLM, where the server utilizes the SLMs’ output logits and cross-entropy losses to enrich the LLM’s understanding.
  5. 5.
    Once the knowledge transfer from SLMs to LLM is completed on the server, the server then computes the LLM’s output logits and the corresponding cross-entropy losses of the public dataset.
  6. 6.
    The server disseminates the LLM’s output logits and the corresponding cross-entropy losses of the public dataset to all the clients.
  7. 7.
  8. 8.
    On the client side, the knowledge transfer flow reverses, with the LLM’s insights being selectively transferred to the SLMs, where the clients leverage the LLM’s output logits and cross-entropy losses to further enhance their SLMs.
3.3 Computation and Communication Complexity
Overall, the computational efficiency and reduced communication complexity of FedMKT make it a promising framework for federated mutual knowledge transfer for large and small Language models, particularly in scenarios where computational resources and communication bandwidth are limited.
@@ 3.3 Computation and Communication Complexity
Table 1: FedMKT: LLM and SLMs Setting
Task Method GPT-2-xlarge OPT-1.3B Bloom-1.3B LLaMa2-1.3B
RTE SLM-ZS 52.4 52.7 52.7 49.8
Standalone 65.7 62.5 58.1 55.6
FedMKT-SLM 70.4 65.7 61.7 58.8
WIC SLM-ZS 49.8 50.8 50 50
Standalone 59.3 52.1 59.1 50.7
FedMKT-SLM 63.2 62.2 61.1 51.9
BoolQ SLM-ZS 61.3 58.4 59.0 61.0
Standalone 71.1 74.1 69.9 69.7
FedMKT-SLM 75.1 76.8 71.4 75.1
CQA SLM-ZS 36.7 41.9 33.8 30.1
Standalone 56.0 58.6 44.7 56.7
FedMKT-SLM 58.3 60.5 50.8 57.0
OpenBookQA SLM-ZS 22.4 23.4 19.6 23.2
Standalone 27.4 26.6 21.0 25.8
FedMKT-SLM 27.8 28.0 23.2 27.2
ARC-E SLM-ZS 58.3 57.0 51.5 53.1
Standalone 59.3 57.9 56.9 60.3
FedMKT-SLM 59.8 59.6 57.5 60.8
ARC-C SLM-ZS 25.0 23.4 23.6 26.7
Standalone 28.2 28.4 24.9 28.5
FedMKT-SLM 30.2 29.4 26.6 30.0
S-NI SLM-ZS 5.0 5.2 5.1 5.8
Standalone 27.9 26.1 10.6 33.4
FedMKT-SLM 34.2 36.0 15.1 37.3
DialogSum SLM-ZS 5.4 6.4 4.9 5.7
Standalone 22.3 19.8 13.1 21.4
FedMKT-SLM 23.2 20.9 14.9 21.6
Table 2: Heterogeneous Setting SLMs Performance Comparison: FedMKT versus Baseline.
@@ 4.1 Setup
We set up a scenario involving three clients and one server to evaluate the FedMKT framework using various LLMs and SLMs.
@@
Models. We evaluate FedMKT on one LLM (LLaMa2-7B Touvron et al. (2023)) in the server, four SLMs in the clients including GPT-2-xlarge (1.5B) Radford et al. (2019), OPT-1.3B Zhang et al. (2022a), Bloom-1.1B Scao et al. (2022) and LLaMa2-1.3B Xia et al. (2023). In our experiments, we evaluate our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. Table 1 details the model configurations for both the LLM and SLMs in the distinct scenarios.
@@
Datasets. We comprehensively evaluate FedMKT on 7 QA datasets and 2 instruction-following datasets. Specifically, for QA tasks, we use RTE Wang et al. (2019), WTC Wang et al. (2019), BoolQ Clark et al. (2019), CommonsenseQA(CQA) Talmor et al. (2018), OpenBookQA Mihaylov et al. (2018), ARC-E Clark et al. (2018), ARC-CClark et al. (2018) to evaluate FedMKT. As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al. (2022), DialogSum Chen et al. (2021).
Baselines. We conduct a comparative analysis of our FedMKT framework against several baselines to evaluate its performance. These baselines included:
  • LLM-ZS, which represents the zero-shot (ZS) capabilities of the large language model (LLM) on the server;
  • SLM-ZS, which represents the zero-shot performance of the small language model (SLM) on the clients;
@@
  • Standalone, where each client independently fine-tunes its local model using its own private dataset;
@@
  • FedAvg, which represents the standard federated averaging algorithm and is only used in the SLM comparisons of homogeneous settings.
@@
  • LLM2SLM, where the LLM is not updated in the FedMKT and is LLM only used to transfer knowledge to SLM in the FedMKT. It is only used in the SLM comparisons of the One-to-One setting.
  • Centralized-LLM, where the server’s LLM is fine-tuned locally using the entirety of the private datasets combined with a public distillation dataset.
Evaluation Metrics. We evaluate the performance of fine-tuned LLM and SLMs on both QA datasets and instruction-following datasets. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L. This approach allows us to comprehensively assess the models’ capabilities across different tasks, providing a clear comparison of their performance.
4.2 Heterogeneous Setting: Performance Evaluations
Tables 2 and 3 demonstrate the performance comparisons of our approach against other baselines in the heterogeneous setting. Specifically, Table 2 focuses on the performance of SLMs, while Table 3 highlights the performance of LLM.
As illustrated in Table 2, FedMKT-SLM demonstrates notably superior performance in comparison to both SLM-ZS and Standalone models. For instance, in the RTE QA dataset, FedMKT-SLM demonstrates significant performance gains compared to both SLM-ZS and Standalone models across various SLMs. Specifically, when using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 34% and Standalone by 7%. Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 25% and Standalone by 5%. In the Bloom-1.3B SLM, FedMKT-SLM achieves a 17% improvement over SLM-ZS and a 6% advantage over Standalone. Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 18% and Standalone by 6%. In the DialogSum instruction-following dataset, FedMKT-SLM again demonstrates impressive results. When using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 330% and Standalone by 4%. Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 227% and Standalone by 6%. In the Bloom-1.3B SLM, FedMKT-SLM achieves a 204% improvement over SLM-ZS and a 14% advantage over Standalone. Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 279% and Standalone by 0.9%. This highlights that within our framework, SLMs are able to effectively leverage the knowledge transferred from LLM, leading to enhanced model capabilities.

### v2

@@ 3.2 Bidirectional Token Alignment
A significant challenge in aligning output logits distributions lies in the mismatch between tokenizers of different LLM and SLMs, exemplified by Bloom and LLaMa. Consider the sentence, "we utilize the dynamic programming approach to align tokens" as an example. Utilizing the Bloom tokenizer would segment it into the following tokens: [’we’, ’utilize’, ’the,’ ’dynamic,’ ’programming,’ ’approach,’ ’to,’ ’align,’ ’tokens’]. However, if the LLaMa tokenizer were used, the segmentation would be: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’].
To tackle this issue, we adopt dynamic programming techniques to promote robust alignment, as evidenced in studies Wan et al. (2024); Fu et al. (2023). Utilizing LLaMa2 and Bloom as illustrative examples, we establish an optimized vocabulary mapping table based on minimum edit distance (MinED). This mapping table identifies the closest Bloom token for each LLaMa2 token (e.g., ’utilize’ for ’util’). We then tokenize a sentence using both tokenizers and apply a dynamic programming algorithm to determine the optimal matching path. When multiple LLaMa2 tokens align to a single Bloom token (e.g., ’util’ and ’ize’ aligning to ’utilize’), we handle them according to the mapping table. Please refer to Appendix B for more details.
In FedMKT, a bidirectional token alignment process occurs before knowledge transfer between LLMs and SLMs. One the one hand, when clients transfer knowledge from their SLMs to the server’s LLM, the server aligns SLM tokens to LLM tokens. On the other hand, when the server transfers knowledge from its LLM back to clients’ SLMs, each client aligns LLM tokens to its SLM tokens.
3.3 Selective Mutual Knowledge Transfer Between LLM and SLMs
To transfer knowledge between the server and clients efficiently, we leverage LoRA to fine-tune the server’s LLM and clients’ SLMs. Specifically, each client $k$ inserts a small low-rank adapter parameterized by $\theta_{k}$ into its local SLM. We denote client $k$ local SLM with the added $\theta_{k}$ as $g_{\phi_{k}+\theta_{k}}$. Likewise, the server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$. We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$. During the whole federated learning training process, $\theta_{k},k=1,...,K$ and $\omega$ are trained, while $\phi_{k},k=1,...,K$ and $\psi$ are frozen.
Before transferring knowledge to the server, each client $k$ trains its LoRA adapter $\theta_{k}$ using its private dataset $\mathcal{D}_{k}$. Consequently, Eq.(1) can be reformulated as follows:
$\mathcal{L}_{1}(\theta_{1},\theta_{2},...,\theta_{K};\{\mathcal{D}_{k}\}_{k=1}^{K})$ (4)
$=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\ell_{\text{TA}}($ $g_{\phi_{k}+\theta_{k}}(x),y)$
@@ 3.3 Selective Mutual Knowledge Transfer Between LLM and SLMs
where $\ell_{\text{TA}}$ is the task loss for training $\theta_{k}$ of each client $k$. The original model parameter $\phi_{k}$ of client $k$’s SLM is frozen during training.
@@
Then, both the server and clients fine-tune their LoRA adapters based on a shared public dataset $\mathcal{D}_{p}$. We formulate the losses of fine-tuning $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
$\mathcal{L}^{f}_{\text{FT}}(\omega;\mathcal{D}_{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{p}}\ell_{\text{CE}}(f_{\psi+\omega}(x),y)$ (5)
$\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}_{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
@@
Next, the server and clients conduct selective knowledge transfer to each other. The motivation for applying selective knowledge transfer is that some clients’ knowledge may adversely affect the performance of LLM on the server and vice versa in a heterogeneous environment. Therefore, it is critical to guarantee that the knowledge transferred between the server and clients is positive to the performance of LLM and SLMs. To this end, we propose a selective knowledge transfer strategy on both the server and client sides, termed DualMinCE.
DualMinCE aims to select knowledge that is positive to the performance of the server’s LLM from clients and vice versa. Specifically, when knowledge needs to be transferred from SLMs to the LLM, each client $k$ computes a knowledge set $\mathcal{S}_{k}=\{l^{i}_{k},p^{i}_{k}\}_{i=1}^{N}$ consisting of loss-logit pairs through its local model based on the public dataset $\mathcal{D}_{p}$. Then, all $K$ clients send their $\{\mathcal{S}_{k}\}_{k=1}^{K}$ to the server. By leveraging DualMinCE (see Algorithm 2 for detail), the server picks a logit $p_{k^{*}}^{i}$ with the smallest loss from $\{l^{i}_{k},p^{i}_{k}\}_{k=1}^{K}$ and adds $p_{k^{*}}^{i}$ to a selective knowledge set $\tilde{\mathcal{S}}_{0}$ if the loss $l_{k^{*}}^{i}$ of $p_{k^{*}}^{i}$ is smaller than the loss $l_{\text{local}}^{i}$ computed through the server’s local LLM based on $x^{i}$ for each $x^{i}$ in $\mathcal{D}_{p}$.
Next, the server leverages the knowledge distillation loss, denoted as $\mathcal{L}_{\text{KD}}^{f}$, to fine-tune $f_{\psi+\omega}$:
$\mathcal{L}^{f}_{\text{KD}}(\omega;\tilde{\mathcal{S}}_{0})=$ $\mathbb{E}_{(x,p)\sim\tilde{\mathcal{S}}_{0}}\ell_{\text{CE}}(f_{\psi+\omega}(x),p)$ (7)
Likewise, each client $k$ leverages DualMinCE to form its selective knowledge set $\tilde{\mathcal{S}}_{k}$ from the knowledge $\mathcal{S}_{0}$ sent from the server. Each client $k$ leverages the following knowledge distillation loss to fine-tune its local model $g_{\phi_{k}+\theta_{k}}$:
$\mathcal{L}^{g}_{\text{KD}}(\theta_{k};\tilde{\mathcal{S}}_{k})=$ $\mathbb{E}_{(x,p)\sim\tilde{\mathcal{S}}_{k}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),p)$ (8)
@@
where $\lambda$ is the hyperparameter that controls the weight of mutual knowledge transfer.
@@
Table 1: The five different settings we utilize to evaluate FedMKT.
Task Method GPT-2-xlarge OPT-1.3B Bloom-1.1B LLaMa2-1.3B LLaMa2-7B
RTE Centralized - - - - 85.9
Zero-Shot 52.4 52.7 52.7 49.8 63.2
Standalone 65.7 62.5 58.1 55.6 -
FedMKT 70.4 65.7 61.7 58.8 82.3
WIC Centralized - - - - 70.4
Zero-Shot 49.8 50.8 50 50 50.3
Standalone 59.3 52.2 59.1 50.6 -
FedMKT 63.2 62.2 61.1 51.9 61.3
BoolQ Centralized - - - - 87.6
Zero-Shot 61.3 58.4 59.0 61.0 70.1
Standalone 71.1 74.1 69.7 69.9 -
FedMKT 75.1 76.8 71.4 75.1 85.0
CQA Centralized - - - - 69.5
Zero-Shot 36.7 41.9 33.8 30.1 39.5
Standalone 56.0 58.6 44.7 56.7 -
FedMKT 58.3 60.5 50.8 57.0 71.8
ARC-E Centralized - - - - 76.9
Zero-Shot 58.3 57.0 51.5 53.1 69.3
Standalone 59.3 57.9 56.9 60.4 -
FedMKT 59.8 59.6 57.5 60.8 76.1
ARC-C Centralized - - - - 48.9
Zero-Shot 25.0 23.4 23.6 26.7 40.0
Standalone 28.2 28.4 24.9 28.5 -
FedMKT 30.2 29.4 26.6 30.0 44.7
S-NI Centralized - - - - 49.3
Zero-Shot 5.0 5.2 5.1 5.8 12.0
Standalone 27.9 26.1 10.6 33.4 -
FedMKT 34.2 36.0 15.1 37.3 41.4
DialogSum Centralized - - - - 27.7
Zero-Shot 5.4 6.4 4.9 5.7 8.5
Standalone 22.3 19.8 13.2 21.4 -
FedMKT 23.2 20.9 14.9 21.6 24.2
Table 2: Method Performance Comparison in the Heterogeneous setting. We evaluate FedMKT with 8 different tasks. In all the 8 tasks, the server is deployed with a LLaMa2-7B model, and the 4 clients are deployed with a GPT-2-xlarge, a OPT-1.3B, a Bloom-1.1B, and a LLaMa2-1.3B, respectively. The ’-’ indicates a method does not apply to the corresponding participant (either the server or the client).
@@ 4.1 Setup
We set up a federated learning scenario involving four clients and one server to evaluate the FedMKT using various publicly available LLMs and SLMs.
@@
Models. We evaluate FedMKT on one LLM (LLaMa2-7B Touvron et al. (2023)) in the server, four SLMs in the clients including GPT-2-xlarge (1.5B) Radford et al. (2019), OPT-1.3B Zhang et al. (2022a), Bloom-1.1B Scao et al. (2022) and LLaMa2-1.3B Xia et al. (2023). In our experiments, we evaluate our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. Table 1 details the setup for the LLM and SLMs in different settings.
@@
Datasets. We evaluate FedMKT comprehensively on 6 QA datasets and 2 instruction-following datasets. Specifically, for QA tasks, we use RTE Wang et al. (2019), WTC Wang et al. (2019), BoolQ Clark et al. (2019), CommonsenseQA(CQA) Talmor et al. (2018), ARC-E Clark et al. (2018), ARC-C Clark et al. (2018) to evaluate FedMKT. As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al. (2022), DialogSum Chen et al. (2021).
Baselines. We conduct a comparative analysis of FedMKT against the following baselines:
  • Centralized, in which the server’s LLM is fine-tuned locally using the datasets combining private datasets of involved clients and the public dataset. In the One-to-One setting, the data of one client and the public data are used to fine-tune the server’s LLM, whereas in other settings, the data of all four clients and the public data are used to fine-tune the LLM;
  • Zero-Shot, representing the zero-shot capabilities of LLM or SLMs (without fine-tuning);
@@
  • Standalone, in which each client independently fine-tunes its local SLM using its private dataset;
@@
  • FedAvg, representing the standard federated averaging algorithm. FedAvg is only used in homogeneous settings because it requires all clients’ models have the same architecture.
@@
  • LLM2SLM, representing FedMKT involving one server with an LLM and one client with an SLM. The LLM is not updated and is used to transfer knowledge to SLM. LLM2SLM is only used in the One-to-One setting.
Evaluation Metrics. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L.
4.2 Evaluation on Heterogeneous Setting
In the Heterogeneous setting, the server is deployed with a LLaMa2-7B model, and the 4 clients are deployed with a GPT-2-xlarge, a OPT-1.3B, a Bloom-1.1B, and a LLaMa2-1.3B, respectively. Table 2 reports the performance comparisons of FedMKT against baselines on 8 tasks.
Tables 2 show that FedMKT performs superior over Zero-Shot and Standalone on all clients’ SLMs. Take the RTE dataset as an example, FedMKT outperforms Zero-Shot by 34% and Standalone by 7% on the GPT-2-xlarge SLM; FedMKT surpasses Zero-Shot by 25% and Standalone by 5% on the OPT-1.3B SLM; FedMKT-SLM achieves a 17% improvement over Zero-Shot and a 6% improvement over Standalone on the Bloom-1.1B SLM; FedMKT-SLM outperforms Zero-Shot by 18% and Standalone by 6% on the LLaMa2-1.3B SLM. These empirical results demonstrate that, by leveraging FedMKT, SLMs are able to effectively leverage the knowledge transferred from the LLM, leading to enhanced model capabilities.